OpenStudy (anonymous):

Solve for x. 0 = 5x^(4/3) - 6x^(2/3) + 1

OpenStudy (anonymous):

Rearrange it: \[\large 5x^{\frac{4}{3}} - 6 x^{\frac{2}{3}} + 1 = 0\]

OpenStudy (anonymous):

\[\large Let \; \; \; x^{\frac{2}{3}} = y\]

OpenStudy (anonymous):

So: \[5y^2 - 6y +1 =0 \]

OpenStudy (anonymous):

Solve for \(y\) now..

OpenStudy (anonymous):

Thank you!